# Why do Learning Rates Transfer? Reconciling Optimization and Scaling Limits for Deep Learning

### v1

@@ Abstract
Recently, there has been growing evidence that if the width and depth of a neural network are scaled toward the so-called rich feature learning limit ($\mu$P and its depth extension), then some hyperparameters — such as the learning rate — exhibit transfer from small to very large models, thus reducing the cost of hyperparameter tuning. From an optimization perspective, this phenomenon is puzzling, as it implies that the loss landscape is remarkably consistent across very different model sizes. In this work, we find empirical evidence that learning rate transfer can be attributed to the fact that under $\mu$P and its depth extension, the largest eigenvalue of the training loss Hessian (i.e. the sharpness) is largely independent of the width and depth of the network for a sustained period of training time. On the other hand, we show that under the neural tangent kernel (NTK) regime, the sharpness exhibits very different dynamics at different scales, thus preventing learning rate transfer. But what causes these differences in the sharpness dynamics? Through a connection between the spectra of the Hessian and the NTK matrix, we argue that the cause lies in the presence (for $\mu$P) or progressive absence (for the NTK regime) of feature learning, which results in a different evolution of the NTK, and thus of the sharpness. We corroborate our claims with a substantial suite of experiments, covering a wide range of datasets and architectures: from ResNets and Vision Transformers trained on benchmark vision datasets to Transformers-based language models trained on WikiText.
${}^{1}$ Zürich, ${}^{2}$ Institute Tübingen, ${}^{3}$ for Intelligent Systems, ${}^{4}$übingen AI Center
@@ 1 Introduction
Figure 1: Top row: learning rate transfer for $\mu$P, but not for NTP. Bottom row: sharpness’ dynamics. Notice how the sharpness is width-independent for $\mu$P, but is decreasing in NTP. Also notice how in this case for $\mu$P the optimal learning rate corresponds to a sharpness that is close to EoS (black dashed lines). We argue that a width-independent sharpness is at the hearth of learning rate transfers. The architecture is a two-layer convolutional network trained on CIFAR-10 with data augmentation, where the width corresponds to the number of filters in the convolution. (See App. B). The width in the legend here refers to the size of the final linear readout layer. Other parameters: $B=128$, epochs $=50$.
@@
While there exist several ways to scale up the width and depth of the network, not all of them facilitate learning rate transfer. For standard deep learning practices, such as networks initialized with LeCun/Kaiming techniques (LeCun, 2002, He et al., 2015), a significant shift in the optimal learning rate is usually observed as the width and the depth of the model are increased. Similarly, under the NTK regime Jacot et al. (2018), which provides theoretical insights into the linearized behavior of very wide neural networks during training, the optimal learning rate also varies as the width and depth of the network change [^1] NTK differs from standard LeCun initialization in that instead of scaling initialization variances, scaling factors are placed as multipliers across the model definition, changing the network parameterization and therefore gradients during backpropagation – without affecting the forward pass distribution at initialization.. Alternatively, Yang and Hu (2021), Yang et al. (2022) propose the $\mu$P framework, designed to maximize the gradient update of the representations of the intermediate layers (i.e. feature learning) as the width increases. Under $\mu$P scaling, and its depth extension for residual networks $\mu P-\sqrt{\text{depth}}$ (Bordelon et al., 2023, Yang et al., 2023), it has been empirically demonstrated that the learning rate transfers across both width and depth.
From an optimization theory viewpoint, however, the rationale behind the connection between the nature of the scaling limit and learning rate transfer remains unclear. In the theory of smooth convex (Nesterov, 2013), nonconvex (Garrigos and Gower, 2023), and stochastic (Gower et al., 2019) optimization, the sharpness — i.e. the maximum Hessian eigenvalue — plays a crucial role in the selection of the optimal step size, and in the guarantees of convergence of gradient methods. As the Hessian dynamics are in general dependent on the network’s width and depth, it is not obvious why the optimal learning rate of a small model should transfer to a larger one.
In this paper, we bridge the gap between the nature of the scaling limit and the loss landscape by comparing the NTK and $\mu$P regimes. We unveil the surprising fact that for $\mu$P and its depth extension, the sharpness remains largely independent of the width and depth of the network. Instead, it progressively converges to a threshold mostly dependent on the batch size and learning rate, which in some cases corresponds to the Edge of Stability (EoS) value (Cohen et al., 2021). Once the sharpness converges to this threshold, it oscillates around it for a sustained period of training time, maintaining dynamics which are largely width/depth-independent. On the other hand, we show that under the NTK regime, the sharpness dynamics significantly separate during training for different widths. Small-width models are still converging to the limiting threshold, while larger width models progressively slow down their convergence. More concretely:
  • We show that under $\mu$P, after a progressive sharpening phase, the dynamics of the largest Hessian eigenvalue $\lambda$ stabilize to a width-independent value (Fig. 1, top left). This correlates with the phenomenon of hyperparameter transfer (Fig. 1, bottom left) as expected by optimization theory. On the other hand, we show in Fig. 1 (right column) that this is not the case for the NTK regime, where the sharpness is decreasing in width. Also, here we do not observe learning rate transfer.
  • Thanks to the depth extension $\mu$P $-\sqrt{\text{depth}}$, we show that our observations remain valid when the depth is scaled up, testing networks with up to $200$ layers.
  • We show that the results on width/depth-independent sharpness are consistent at realistic scale, including ResNets and Vision Transformers (ViTs) trained on Imagenet and GPT-2 on text data.
  • We investigate the threshold that is reached during the progressive sharpening phase (in $\mu$P), connecting our findings to the line of work of Edge of Stability Cohen et al. (2021). In particular, EoS is achieved width/depth-independently for large batch sizes and relatively large learning rates, while for small batches the sharpness stays below EoS (while still being largely independent of the width/depth).
These results on width/depth-independent sharpness provide a clean interpretation for learning rate transfer: under $\mu$P, the key landscape features along the trajectory of gradient descent are remarkably similar across different scales. But why does this hold only for $\mu$P but not for the neural tangent parameterization (NTP) or other initialization/parameterization strategies? We argue that while in $\mu$P feature learning guarantees progressive sharpening to reach a width-independent sharpness at any scale, in NTP the progressive lack of feature learning when the width is increased prevents the Hessian from adapting, and its largest eigenvalue from reaching the convergence threshold. More concretely, in Section 3.1 we show that the progressive sharpening phase is mainly driven by the evolution of the NTK’s largest eigenvalue, which is asymptotically fixed to its initial value for NTP, while it evolves at any width under $\mu$P. This disparity elucidates the distinct behaviors of sharpness under these parameterizations, shedding light on the intricate interplay between scaling limits and optimization dynamics. We hope that our results help reduce conceptual gaps between the lines of research on scaling limits and optimization.
The rest of the paper is organized as follows. In Section 2 we provide the necessary background on scaling limits, optimization of neural networks, and EoS. In Section 3 we present the main empirical results on the connection between sharpness and hyperparameter transfer while in Section 4 we provide further intuition with a theoretical analysis on a two-layer linear network. Finally, in Section 5 we discuss the relevance of our results in the existing literature and conclude with potential future directions. We defer a more extensive discussion on related work to the appendix (Sec. A).
2 Background
2.1 Scaling Limits of Neural Networks
@@
$h^{\ell+1}(x)=h^{\ell}(x)+\frac{1}{\sqrt{N}L^{\alpha}}W^{\ell}\phi(h^{\ell}(x)),$ (1)
where $N$ and $L$ are the width and depth of the network, $W^{\ell}\in\mathbb{R}^{N\times N}$ for $\ell=1,\dots,L-1$. We denote the output with $f(x)=\frac{1}{\gamma}W^{L}\phi(h^{L}(x))$, where $W^{L}\in\mathbb{R}^{1\times N}$ and $\gamma$ scales the network output and has a crucial role on transfer. Similarly, $\alpha$ has the role of interpolating between different depth limit regimes. At the first layer, we define $h^{1}(x)=\frac{1}{\sqrt{D}}W^{0}x$, where $W^{0}\in\mathbb{R}^{N\times D}$. All the weights $\theta=\{W^{\ell}\}_{l=0}^{L}$ are initialized independently from $\mathcal{N}(0,1)$ and we denote with $P$ the total number of parameters. We stress that the fully connected layer can be replaced with any type of layer (our experiments include convolutional and attention layers). Given a dataset $\mathcal{D}=\{(x_{\mu},y_{\mu})\}_{\mu=1}^{|\mathcal{D}|}$ of datapoints $x_{\mu}\in\mathbb{R}^{D}$ and labels $y_{\mu}\in\mathbb{R}$, we train the network with stochastic gradient descent (SGD) with batch size $B$ and learning rate $\eta\in\mathbb{R}$:
$\theta_{t+1}=\theta_{t}-\eta\sum_{\mu=1}^{B}\nabla_{\theta}\mathcal{L}(f_{t}(x_{\mu}))$ (2)
where $\mathcal{L}$ is a twice differentiable loss function. Defining $f_{t}:=(f_{t}(x_{\mu}))_{\mu\in[|\mathcal{D}|]}\in\mathbb{R}^{|\mathcal{D}|}$ to be the vector of network’s outputs at time $t$, if one considers continuous time, the corresponding gradient descent dynamics in function space $df_{t}/dt$ take the following form Arora et al. (2019):
$\frac{df_{t}}{dt}=-\Theta(f_{t})\Delta(f_{t}),$ (3)
where $\Delta(f_{t})_{i}:=\partial{\mathcal{L}(f_{t}(x_{i}))}/\partial f_{t}(x_{i})$, $i\in[|\mathcal{D}|]$ is the vector of residuals, and
$\Theta(f_{t})_{ij}:=\langle\nabla_{\theta}f_{t}(x_{i}),\nabla_{\theta}f_{t}(x_{j})\rangle$ (4)
for $i,j\in[|\mathcal{D}|]$ is the neural tangent kernel (NTK).
@@
The parameters $\gamma,\alpha,\eta\in\mathbb{R}$ determine the nature of the scaling limit. If $\gamma=\gamma_{0},\eta=\eta_{0}$ are $\mathcal{O}(1)$ constants with respect to $N,L$ (neural tangent parameterization, or NTP), then the network enters the NTK regime Jacot et al. (2018). Here, in the limit of infinite width, the NTK remains constant to its value at initialization throughout training, i.e. $\Theta(f_{t})=\Theta(f_{0})$ for all $t\geq 0$. Thus, the network’s dynamics become equivalent to a linear model trained on the first order term of the Taylor expansion of the model at initialization Lee et al. (2019). The fact that the NTK is fixed to its value at initialization is associated with the lack of feature learning of the model in the large width limit.
If $\gamma=\gamma_{0}\sqrt{N}$, and $\eta=\eta_{0}\gamma^{2}$ ($\mu$P, or mean-field parameterization), the features evolve in the limit (i.e. the NTK $\Theta(f_{t})$ evolves), and the richer model’s dynamics can be described using either Tensor Programs Yang and Hu (2021) or dynamical mean field theory Bordelon and Pehlevan (2022). Under $\mu$P, Yang et al. (2022) show that hyperparameters transfer across width, in contrast to kernel limits, which we reproduce for our residual network in Fig. 1.
@@
If on top of the $\mu$P framework, the residual branches are scaled with $\alpha=1/2$ ($\mu$P-$\sqrt{L}$ parameterization), then Bordelon et al. (2023) and Yang et al. (2023) show that the infinite width dynamics also admit a feature-learning infinite depth limit. Under $\mu$P-$\sqrt{L}$, the learning rate transfers with both width and depth. In this paper, we compare NTP and $\mu$P regimes as the width is increased, and show that our results extend to depth-scaling using the $\mu$P-$\sqrt{L}$ model.
() Loss curves
() Sharpness dynamics
() Sharpness vs width
2.2 Conditions for hyperparameter transfer

### v2

# Super Consistency of Neural Network Landscapes and Learning Rate Transfer

@@ Abstract
Recently, there has been growing evidence that if the width and depth of a neural network are scaled toward the so-called rich feature learning limit ($\mu$P and its depth extension), then some hyperparameters — such as the learning rate — exhibit transfer from small to very large models. From an optimization perspective, this phenomenon is puzzling, as it implies that the loss landscape is consistently similar across very different model sizes. In this work, we study the landscape through the lens of the loss Hessian, with a focus on its largest eigenvalue (i.e. the sharpness), and find that certain spectral properties under $\mu$P are largely independent of the size of the network, and remain consistent as training progresses. We name this property Super Consistency of the landscape. On the other hand, we show that in the Neural Tangent Kernel (NTK) and other scaling regimes, the sharpness exhibits very different dynamics at different scales. But what causes these differences in the sharpness dynamics? Through a connection between the Hessian’s and the NTK’s spectrum, we argue that the cause lies in the presence (for $\mu$P) or progressive absence (for the NTK scaling) of feature learning. We corroborate our claims with a substantial suite of experiments, covering a wide range of datasets and architectures: from ResNets and Vision Transformers trained on benchmark vision datasets to Transformers-based language models trained on WikiText.
: Equal contribution. Correspondence to: lorenzo.noci@inf.ethz.ch, ameterez@fas.harvard.edu1ETH Zürich, 2ELLIS Tübingen, 3MPI for Intelligent Systems, 4Tübingen AI Center, 5Harvard UniversityThis is a revised version of the paper previously titled “Why do Learning Rates Transfer? Reconciling Optimization and Scaling Limits for Deep Learning”.
@@ 1 Introduction
Figure 1: Top row. Under $\mu$P, (left) the sharpness dynamics are largely identical for the whole training dynamics across different widths, phenomenon that we call Super Consistency. The dashed horizontal lines are the Edge of Stability thresholds. Center: The loss dynamics are similar early in training, but accumulate finite-size effects over time, thus violating Super Consistency. Right: The learning rate transfers from small to large model, suggesting that the loss landscape is Super Consistent across different model sizes. Bottom row. Under NTK parameterization (NTP), the sharpness dynamics show large discrepancies. Also, the learning rate does not transfer. The architecture is a two-layer convolutional network trained on CIFAR-10 with data augmentation, where the width corresponds to the number of filters in the convolution. (See App. J). Other parameters: $B=128$, epochs $=50$.
@@
While there exist several ways (a.k.a parametrizations) to scale up the width and depth of the network, not all of them facilitate learning rate transfer. For standard deep learning practices, such as networks parametrized with LeCun/Kaiming initializations (LeCun et al., 2002, He et al., 2015), a significant shift in the optimal learning rate is usually observed as the width and the depth of the model are increased. Similarly, under the Neural Tangent Kernel (NTK) parametrization Jacot et al. (2018), which provides theoretical insights into the behavior of very wide neural networks during training, the optimal learning rate also varies as the width and depth of the network change. Alternatively, Yang and Hu (2021), Yang et al. (2022) propose the $\mu$P framework, designed to maximize the gradient update of the representations of the intermediate layers (i.e. feature learning) as the width increases. Under $\mu$P scaling, and its depth extension for residual networks Depth-$\mu$P Bordelon et al. (2023), Yang et al. (2023), it has been empirically demonstrated that the learning rate transfers across both width and depth. In Vyas et al. (2023) it is observed that in feature learning parametrizations (e.g. $\mu$P) the model’s dynamics are consistent across model sizes, but for harder tasks or longer training times there are progressive and significant deviations across different model sizes. We give an example in Figure 1 (top center), where the training losses exhibits an increasing gap. The fact that the learning rate is exactly preserved, however, suggests that some properties of the landscape do not exhibit these finite-size deviations, and must be precisely preserved across different model sizes for the whole training trajectory.
Motivated by this, in the present work we identify the notion of Super Consistency to describe the properties of the neural network’s loss landscape that are preserved across training as a function of the model width and depth, thus not accumulating finite-size effects. In particular, we analyze the landscape through the lens of the loss Hessian. It provides insights into the landscape’s local curvature, and its structure for neural networks has been studied in several works (Sagun et al., 2016; 2017, Martens, 2020, Singh et al., 2021, Orvieto et al., 2021). Of great interest in optimization theory is the sharpness, i.e. the top Hessian eigenvalue, which for neural networks exhibit a rapid increase (progressive sharpening) towards a threshold called Edge of Stability (EoS) Lewkowycz et al. (2020), Cohen et al. (2021). However, although a few works have provided early insights Sagun et al. (2016), Cohen et al. (2021), Vyas et al. (2023), the scaling properties of the sharpness and Hessian’s dynamics under different scaling limits remain unexplored. In this work we first present evidence of Super Consistency in the Hessian’s largest eigenvalues, which have been shown to control the curvature along the optimization subspace Gur-Ari et al. (2018). We then focus on the sharpness dynamics, and find that the presence (resp. absence) of Super Consistency correlates well with presence (resp. absence) of learning rate transfer under $\mu$P, NTP and other scaling limits. These results suggest that learning rate transfer happens in super consistent landscapes, as the geometry of the landscape does not significantly change with the network’s size.
Then, we investigate the role of feature learning in the progressive sharpening phase, and argue that while in $\mu$P feature learning causes progressive sharpening to reach a width-independent sharpness, in the NTK regime the progressive lack of feature learning when the width is increased prevents the Hessian from adapting, and its largest eigenvalue from reaching the convergence threshold.
More concretely:
  • We define Super Consistency, and show that under $\mu$P and Depth-$\mu$P it holds for the largest eigenvalues of the loss Hessian (Fig. 2), which converge to a largely width-independent threshold and remains there for the rest of training. On the other hand, we show that other quantities, such as the training loss and the NTK eigenvalues accumulate significant finite-size effects. We quantify the rate of divergence of these quantities with power law fits (Fig. 3).
  • We analyze the relationship between Super Consistency of the sharpness and learning rate transfer across $\mu$P, Depth-$\mu$P, NTP and other parametrizations (Fig. 1, Fig. 4 and Sec. B). For $\mu$P and Depth-$\mu$P, which do transfer, the sharpness stays super consistent, stabilizing to a threshold (Fig. 1, top left), which in some cases corresponds Edge of Stability (Cohen et al., 2021), and oscillates around it for a sustained period of training time. On the other hand, under NTP, Standard Parametrization (SP), or Depth-$\mu$P with multiple layers per residual block, the sharpness dynamics significantly separate during training for different widths, albeit in different ways. Also, here we do not observe transfer.
  • We reproduce some of these results at realistic scale, including ResNets and Vision Transformers (ViTs) trained on Imagenet and GPT-2 on text data. Also, we analyze the effect of batch size, learning rate warm-up, and long training times.
  • In Sec. 4.1 we show that the progressive sharpening phase is mainly driven by the NTK’s largest eigenvalue, which is asymptotically fixed to its initial value for NTP, while it evolves at any width under $\mu$P. In Sec. 5 we provide intuition with a theoretical analysis on a two-layer linear network.
Finally, in Sec. 6 we discuss to what extent Super Consistency of these properties explains learning rate transfer, and the relevance of our results in the existing literature on optimization and scaling limits. Due to page limitations, we defer the discussion on related work to the appendix (App. A).
2 Background and Definitions
@@
$h^{\ell+1}(x)=\tau h^{\ell}(x)+\frac{1}{\sqrt{N}L^{\alpha}}W^{\ell}\phi(h^{\ell}(x)),$ (1)
where $N$ and $L$ are the width and depth of the network, $W^{\ell}\in\mathbb{R}^{N\times N}$ for $\ell=1,\dots,L-1$, and $\tau$ is a factor that enables ($\tau=1$) or disables ($\tau=0$) the skip branch. We denote the output with $f(x)=\frac{1}{\gamma}W^{L}\phi(h^{L}(x))$, where $W^{L}\in\mathbb{R}^{1\times N}$ and $\gamma$ scales the network output. Similarly, $\alpha$ has the role of interpolating between different depth limit regimes. At the first layer, we define $h^{1}(x)=\frac{1}{\sqrt{D}}W^{0}x$, where $W^{0}\in\mathbb{R}^{N\times D}$. All the weights $\theta=\{W^{\ell}\}_{l=0}^{L}$ are initialized independently from $\mathcal{N}(0,1)$ and we denote with $P$ the total number of parameters. We stress that the fully connected layer can be replaced with any type of layer (our experiments include convolutional and attention layers). Given a dataset $\mathcal{D}=\{(x_{\mu},y_{\mu})\}_{\mu=1}^{|\mathcal{D}|}$ of datapoints $x_{\mu}\in\mathbb{R}^{D}$ and labels $y_{\mu}\in\mathbb{R}$, we train the network with stochastic gradient descent (SGD) with batch size $B$ and learning rate $\eta\in\mathbb{R}$,
$\theta_{t+1}=\theta_{t}-\eta\sum_{\mu=1}^{B}\nabla_{\theta}\mathcal{L}(f_{t}(x_{\mu})),$ (2)
where the loss $\mathcal{L}$ is a twice differentiable loss function. Defining $f_{t}:=(f_{t}(x_{\mu}))_{\mu\in[|\mathcal{D}|]}\in\mathbb{R}^{|\mathcal{D}|}$ to be the vector of network’s outputs at time $t$, if one considers continuous time, the corresponding gradient descent dynamics in function space $df_{t}/dt$ take the following form Arora et al. (2019): $\frac{df_{t}}{dt}=-\Theta(f_{t})\Delta(f_{t})$, where $\Delta(f_{t})_{i}:=\partial{\mathcal{L}(f_{t}(x_{i}))}/\partial f_{t}(x_{i})$, $i\in[|\mathcal{D}|]$ is the vector of residuals, and $\Theta(f_{t})_{ij}:=\langle\nabla_{\theta}f_{t}(x_{i}),\nabla_{\theta}f_{t}(x_{j})\rangle$ for $i,j\in[|\mathcal{D}|]$ is the Neural Tangent Kernel (NTK).
@@
The parameters $\gamma,\alpha,\eta\in\mathbb{R}$ determine the nature of the scaling limit. If $\gamma=\gamma_{0},\eta=\eta_{0}$ are $\mathcal{O}(1)$ constants with respect to $N,L$ (neural tangent parameterization, or NTP), then the network enters the NTK regime Jacot et al. (2018). Here, in the limit of infinite width, the NTK remains constant to its value at initialization throughout training, i.e. $\Theta(f_{t})=\Theta(f_{0})$ for all $t\geq 0$. Thus, the network’s dynamics become equivalent to a linear model trained on the first order term of the Taylor expansion of the model at initialization Lee et al. (2019). The fact that the NTK is fixed to its value at initialization is associated with the lack of feature learning of the model in the large width limit. If $\gamma=\gamma_{0}\sqrt{N}$, and $\eta=\eta_{0}\gamma^{2}$ ($\mu$P, or mean-field parameterization), the features evolve in the limit (i.e. the NTK $\Theta(f_{t})$ evolves), and the richer model’s dynamics can be described using either Tensor Programs Yang and Hu (2021) or dynamical mean field theory Bordelon and Pehlevan (2022). Under $\mu$P, Yang et al. (2022) show that the learning rate $\eta_{0}$ as well as other hyperparameters transfer across width, in contrast to kernel limits, which we reproduce for our residual network in Fig. 1.
@@
If on top of the $\mu$P framework, the residual branches are scaled with $\alpha=1/2$ (Depth-$\mu$P), then Bordelon et al. (2023) and Yang et al. (2023) show that the infinite width dynamics also admit a feature-learning infinite depth limit. Under Depth-$\mu$P, the learning rate $\eta_{0}$ transfers with both width and depth. In this paper, we compare NTP and $\mu$P regimes as the width is increased, and show that our results extend to depth-scaling using the Depth-$\mu$P model. We summarize the feature learning parametrizations and report Depth-$\mu$P for Adam in Appendix K.
(a) Largest Hessian eigenvalues - $\mu$P
(b) Largest NTK eigenvalues - $\mu$P
Figure 2: (a) The top Hessian eigenvalues exhibit a progressive increase to a threshold, with larger eigenvalues showing precise Super Consistency, while lower eigenvalues show finite-size accumulation at small width in the initial phase of training. (b) Top eigenvalues of the NTK matrix $\Theta$. As opposed to the top eigenvalues of the Hessian, these exhibit evident finite-size accumulation during training. Model: 3-layer ConvNet, $\tau=0$, $\eta_{0}=0.7$ (optimal). Details in Sec. J.
3 Super Consistency of the Optimization Landscape
In this work, we analyze the landscape through the lens of the preconditioned Hessian $\gamma^{2}H_{t}$, where $H_{t}:=\nabla^{2}_{\theta}\mathcal{L}(\theta_{t}):=\sum_{\mu}\nabla^{2}_{\theta}\mathcal{L}(f_{t}(x_{\mu}))\in\mathbb{R}^{P\times P}$, as $\theta_{t}$ evolves with gradient descent. The Hessian is a key object in optimization theory (Nesterov, 2013), information geometry (Amari, 1997; 1998), and deep learning theory (Gur-Ari et al., 2018, Ghorbani et al., 2019, Cohen et al., 2021, Singh et al., 2021) and its relation to optimal step sizes is often used to design second-order optimizers (Amari, 1998, Martens, 2016, Martens and Grosse, 2015, Botev et al., 2017). In Figure 1, we can observe that learning rate transfer correlates with strong alignment across model sizes of the Hessian top eigenvalue dynamics, a property which we name Super Consistency. The choice of the preconditioning factor $\gamma^{2}$ ensures the right scaling with respect to the width $N$, as the theory will justify. We also provide an intuition and an extension to Adam in Appendix. J.1. Unless stated otherwise, every experiment is conducted with the this preconditioning factor $\gamma^{2}$ set according to the corresponding parametrization.
More concretely, Super Consistency refers to when certain aspects of the loss landscape and of the predictor $S_{N}(t)$ (in this paper $S_{N}(t)$ refers to the NTK’s and loss Hessian’s eigenvalues or the loss itself) exhibit the following two properties:
  • At realistically large $N$, $S_{N}(t)$ does not deviate significantly from its limit $S_{\infty}(t):=\lim_{N\to\infty}S_{N}(t)$. This is what is referred to as consistency in Vyas et al. (2023).
  • $S_{N}(t)$ does not accumulate significant finite-size effects over time, i.e. the curves of $S_{N}(t)$ and $S_{\infty}(t)$ remain close over a sustained period of training.
With respect to the experiment illustrated in Fig. 1, notice that the curves of the loss (center) at different widths show progressive and significant deviations, thus violating Super Consistency. On the other hand, the sharpness dynamics for $\mu$P qualitatively exhibit little-to-no deviations. Also, notice that we assume existence of the limit $\lim_{N\to\infty}S_{N}(t)$. For those parametrizations (e.g. standard parametrizationYang et al. (2022)) that do not have a well-defined limit, $S_{N}(t)$ diverges at large $N$ and Super Consistency is trivially violated.
We now turn to the analysis of the Hessian spectrum and observe the following:
Observation: in $\mu$P (and Depth-$\mu$P), Hessian eigenvalues are super consistent along the optimisation trajectory. Smaller eigenvalues have progressively different dynamics.
In Fig. 2 (a), we train a residual network on CIFAR-10 (a 10 classes image classification task) using cross-entropy loss, and show super consistent dynamics of three of the top $k=10$ eigenvalues. The choice of $k=10$ is guided by Gur-Ari et al. (2018), where it is observed that stochastic gradient descent happens in a small subspace where the gradient lies in the space spanned by top $k$ Hessian eigenvectors (where $k$ is the number of classes). Thus, our results show that the curvature along the training trajectory is preserved super consistently at different scales, thus suggesting that the geometry of the landscape across the trajectory is preserved across model size. Lower order eigenvalues tend to accumulate finite-size effects in the first phase of training, and stabilize at lower thresholds for smaller width models. We discuss the effect of lower order eigenvalues through the Hessian trace in Appendix G. Finally, to make sure that Super Consistency holds along the training trajectory regardless of the tiny-subspace assumption, in Appendix I we track the directional sharpness, that measures the curvature along the gradient direction.
To give a quantitative measure to the finite-size accumulation property, we measure deviations over time by estimating the following quantity:
$g(t):=|S_{N}(t)-S_{\infty}(t)|.$ (3)
When $g(t)$ increases over time (up to fluctuations), Super Consistency is violated. We illustrate this in Fig. 3 (b, c), where we compute the left hand side of Eq. 3 for the loss $\mathcal{L}(f_{t})$ and the NTK’s largest eigenvalue $\lambda_{max}(\Theta)$. To estimate the infinite width limit, we use a very large-width model as a proxy. Notice how under $\mu$P the the loss dynamics progressively diverge from the infinite width model, indicating a finite-size accumulation over time. The same holds for $\lambda_{max}(\Theta)$. To study the rate of divergence $g(t)$, we fit a power law of the form $y=at^{\beta}$ to the observations. A larger $\beta$ indicates a higher divergence rate. Notice how $\beta>0.6$ for the loss, and $\beta\approx 2$ for $\lambda_{max}(\Theta)$, indicating quadratic divergence. In comparison, in Fig. 3 (left), we show Super Consistency of the sharpness, in that finite-size effects are not accumulated over time (10 epochs). Finally, both in Fig. 2 and 3 notice how the sharpness is not just converging in width, but in fact width-independent. This might be due to the fact that the threshold is a stable attractor of the dynamics Kalra et al. (2023).
(a) Sharpness vs Time - $\mu$P
(b) Loss vs Time - $\mu$P
(c) NTK $\lambda_{\text{max}}(\Theta)$ vs Time - $\mu$P
Figure 3: (a) Convergence rate of the sharpness at finite width $N$ to the infinite limit proxy. Note that the distance approaches 00 as the training time increases. (b) Convergence rate of the loss at finite width $N$ to the infinite limit proxy. Note that the loss accumulates finite-size effects over time and the distance to the proxy increases. (c) Convergence rate of the top NTK eigenvalues over time to the infinite limit proxy. Similar to the loss, this also accumulates finite-size effects over time. Details: infinite limit proxy is width $4096$, model is ConvNet, $\tau=0$, $\eta_{0}=0.7$.
4 Super Consistency and Learning Rate Transfer
Sharpness and Edge of Stability. We now focus on the sharpness $\lambda:=\lambda_{\max}(\gamma^{2}H)$, defined as the largest eigenvalue of the Hessian. In the theory of smooth convex (Nesterov, 2013), nonconvex (Garrigos and Gower, 2023), and stochastic (Gower et al., 2019) optimization, the sharpness plays a crucial role in in the guarantees of convergence of gradient methods and selection of the optimal step size. For instance, for a quadratic objective, $\lambda_{t}=\lambda$ is constant and gradient descent would diverge if the learning rate satisfies $\eta_{0}>2/\lambda$, and training speed is maximized for $\eta_{0}=1/\lambda$ (LeCun et al. (2002), page 28). Beyond this classical example, which assumes constant Hessian, the descent lemma (Nesterov, 2013) states that $\mathcal{L}(\theta_{t+1})\leq\mathcal{L}(\theta_{t})$ if $\eta\leq\frac{2}{\beta}$ where $\beta:=\sup_{\theta}\|\nabla^{2}L(\theta)\|_{2}$, and $\|\nabla^{2}L(\theta)\|_{2}$ is the sharpness at $\theta$. When it comes to deep neural networks, $\lambda_{t}$ is generally observed to increase during training (progressive sharpening): in the early phase of training it increases Jastrzkebski et al. (2018), Jastrzebski et al. (2020) and then it decreases close to convergence Sagun et al. (2016). Under full batch gradient descent training, the sharpness consistently rises above the EoS threshold of $2/\eta_{0}$ Cohen et al. (2021).